D14610:

A ramp into a building forms a 6° angle with the ground. If the ramp is 8 feet long, how far away from the building is the entry point of the ramp? Round the solution to the nearest hundredth. 5.33 feet 6.05 feet 7.04 feet 7.96 feet

AnnahTheGremlin:

Its it 7.96 feet..? Im not good at math ngl

D14610:

@annahthegremlin wrote:
Its it 7.96 feet..? Im not good at math ngl
how sure are you

H1m:

@d14610 wrote:
@annahthegremlin wrote:
Its it 7.96 feet..? Im not good at math ngl
how sure are you
u still need help?

D14610:

@h1m wrote: @d14610 wrote: @annahthegremlin wrote: Its it 7.96 feet..? Im not good at math ngl how sure are you u still need help? yeahh bro i got like 6 more questions and if i dont pass this exam my parents are gonna beat my Chicken can i dm u?

H1m:

@d14610 wrote:
@h1m wrote: @d14610 wrote: @annahthegremlin wrote: Its it 7.96 feet..? Im not good at math ngl how sure are you u still need help? yeahh bro i got like 6 more questions and if i dont pass this exam my parents are gonna beat my Chicken can i dm u?
Gladly always!

grimreaper124:

idk sorry

AnnahTheGremlin:

Not trying to blame anyone but when you said "if i don't pass this exam my parents are gonna beat my" y'know, i sorta got in trouble at my school, I'm just saying respectfully.

surjithayer:

let the distance be x. \[\frac{ x }{ 8 }=\cos 6,x=8\cos 6\approx7.956~ft\] \approx ~7.96 `ft

surjithayer:

7.96 ft

Twaylor:

bros cheating on homework xD

pgray:

of course

pgray:

they stupitie

pgray:

stupit